OpenStudy (anonymous):

how do you solve 4/15-3/5=x-4/3

hero (hero):

Multiply both sides by 15 to eliminate the denominators.

OpenStudy (anonymous):

how do you muliply both sides by 15

hero (hero):

\(15\left(\dfrac{4}{15} - \dfrac{3}{5}\right) = 15\left(x - \dfrac{4}{3}\right)\)

OpenStudy (anonymous):

do i need to remove the parentheses

hero (hero):

You can only remove the parentheses after apply the distributive rule.

hero (hero):

\(15\left(\dfrac{4}{15}\right) - 15\left(\dfrac{3}{5}\right) = 15x - 15\left(\dfrac{4}{3}\right)\)

OpenStudy (anonymous):

still confused.

OpenStudy (texaschic101):

by multiplying both sides by the common denominator...this gets rid of the fractions

OpenStudy (anonymous):

15x=4/3

OpenStudy (texaschic101):

or you can do it with the fractions. 4/15 - 3/5 = x - 4/3 --- add 4/3 to both sides 4/15 - 3/5 + 4/3 = x now...to add/subtract fractions you need a common denominator. A number that all your denominators go into evenly. That would be 15. 4/15 = 4/15 -3/5 = - 9/15 4/3 = 20/15 now we re-write the problem 4/15 - 9/15 + 20/15 = x now do the math and what do you get ?

OpenStudy (anonymous):

43/15

OpenStudy (texaschic101):

(4 - 9 + 20) / 15 = x 15/15 = x 1 = x check.. 4/15 - 3/5 = x - 4/3 4/15 - 9/15 = 1 - 20/15 4/15 - 9/15 = 15/15 - 20/15 -5/15 = -5/15 (correct) so x = 1

OpenStudy (anonymous):

oh ok I see. Thank You

OpenStudy (texaschic101):

YW :)